OpenStudy (anonymous):

nina and philip cut mr fergerson lawn nina worked alone on her half but philip shared his half equally with his friends ezera and ben what fraction of the earnings should each person get

OpenStudy (anonymous):

Nina should get 1/2 because she did half the yard. Philip, Ezera, and Ben should EACH get 1/3 of 1/2 = 1/3 x 1/2 = 1/6